Can someone help me on this?
I need to divide:
3 3/4 divided by 1 3/8
Thanks,
Lyn

Found 2 solutions by tutorcecilia, stanbon:
Answer by tutorcecilia(2152)

3 3/4 divided by 1 3/8 [Change each fraction to an improper fraction]
(15/4) divided by [11/8]
(15/4) x (8/11) [When dividing by fractions, flip the denominator and multiply]
(15/1) x (2/11) [Cancel wherever possible and multiply straight across]
30/11 {Change to an improper fraction]
2 8/11 [Final answer]

Answer by stanbon(75887)

3 3/4 divided by 1 3/8
3 3/4 = 15/4
1 3/8 = 11/8
[15/4]/[11/8] = [15/4]
Reduce the 8 wit the 4 to get:
[15/1]
Multiply to get:
[30/11]
=2 8/11
Cheers,
Stan H.